using perimeter formula, find the perimeterof the retangle in terms of x.
top is 3x-8
side is x

Answer by jim_thompson5910(35256)

L = 3x - 8
W = x

--------------------------------------------------------------

P = 2*(L + W)


P = 2*(3x-8 + x)


P = 2*(4x-8)


P = 2*(4x)+2(-8)


P = 8x - 16